OpenStudy (anonymous):

log sub a5=.699 and log sub a 3=.477. Use these values to evaluate log sub a 45.

OpenStudy (mathstudent55):

Is this your question? \( \log_a 5 = 0.699\) \( \log_a 3 = 0.477\) Find \(\log_a 45\)

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

Remember these two important rules of logs: \(\log_a x^m = m \log_a x \) \( \log_a (xy) = \log_a x + \log_a y\)

OpenStudy (mathstudent55):

Can you break down 45 into the product of its prime factors?

OpenStudy (anonymous):

how were you able to type the sub a?

OpenStudy (mathstudent55):

I'm using the equation editor.

OpenStudy (anonymous):

I came up with 1.653 as my answer but not sure if it's right. i did 2(.477) +.699

OpenStudy (anonymous):

because 3^2 * 5 would equal 45

OpenStudy (mathstudent55):

That's it. Here's how you do it. \(45 = 3^2 \cdot 5\) \(\log_a 45\) \( = \log_a (3^2 \cdot 5)\) \( = \log_a 3^2 + \log_a 5 \) \(= 2 \log_a 3 + \log_a 5\) \(= 2 \cdot 0.477 + 0.699\)

OpenStudy (mathstudent55):

Which is what you did. Great job!

OpenStudy (anonymous):

thank you so much!

OpenStudy (mathstudent55):

You're very welcome.